# Curated LLM: Synergy of LLMs and Data Curation for tabular augmentation in ultra low-data regimes

### v1

@@ C.2 Full results for performance evaluation
seer (n=40) $87.90_{(0.09)}$ $82.93_{(0.55)}$ $84.29_{(0.39)}$ $\textbf{84.93}_{(0.46)}$ $83.46_{(1.20)}$ $84.44_{(0.69)}$ $80.05_{(0.93)}$ $83.67_{(0.51)}$ $82.59_{(1.48)}$ $81.37_{(1.11)}$ $54.93_{(2.06)}$ $81.11_{(1.24)}$ $79.88_{(0.53)}$ $84.36_{(0.62)}$ $80.79_{(0.67)}$ $82.21_{(0.67)}$ $78.69_{(2.32)}$ $83.62_{(1.06)}$
compas (n=40) $67.35_{(0.16)}$ $62.34_{(0.79)}$ $67.57_{(0.40)}$ $\textbf{67.85}_{(0.37)}$ $61.34_{(1.67)}$ $62.84_{(1.21)}$ $56.29_{(1.96)}$ $61.02_{(1.69)}$ $58.85_{(1.28)}$ $60.11_{(1.22)}$ $58.88_{(1.04)}$ $64.37_{(0.91)}$ $58.61_{(1.35)}$ $63.54_{(1.11)}$ $60.83_{(1.19)}$ $60.95_{(1.22)}$ $55.94_{(1.50)}$ $61.04_{(1.56)}$
adult (n=40) $84.43_{(0.06)}$ $79.44_{(1.03)}$ $48.31_{(3.58)}$ $73.82_{(2.19)}$ $49.21_{(1.78)}$ $74.27_{(1.49)}$ $71.82_{(1.22)}$ $79.11_{(0.71)}$ $71.51_{(1.85)}$ $77.99_{(0.65)}$ $66.77_{(3.32)}$ $78.81_{(1.36)}$ $71.13_{(2.27)}$ $79.71_{(1.02)}$ $77.90_{(1.07)}$ $78.84_{(1.19)}$ $72.58_{(1.34)}$ $\textbf{80.02}_{(0.76)}$
drug (n=40) $77.71_{(0.32)}$ $71.86_{(1.07)}$ $74.30_{(0.59)}$ $\textbf{75.79}_{(0.39)}$ $71.33_{(0.88)}$ $72.76_{(0.97)}$ $69.46_{(2.23)}$ $72.74_{(1.70)}$ $71.08_{(1.72)}$ $73.07_{(1.03)}$ $64.89_{(1.39)}$ $73.64_{(0.87)}$ $62.51_{(3.03)}$ $70.97_{(1.91)}$ $69.23_{(1.68)}$ $69.78_{(1.46)}$ $65.22_{(1.37)}$ $70.30_{(1.04)}$
covid (n=100) $74.52_{(0.16)}$ $71.57_{(0.48)}$ $73.77_{(0.27)}$ $\textbf{74.71}_{(0.34)}$ $70.71_{(0.46)}$ $72.76_{(0.44)}$ $69.05_{(0.96)}$ $72.13_{(0.66)}$ $71.60_{(0.59)}$ $73.22_{(0.46)}$ $63.52_{(1.29)}$ $72.04_{(0.57)}$ $64.25_{(1.51)}$ $72.64_{(0.64)}$ $70.08_{(0.67)}$ $70.78_{(0.59)}$ $69.05_{(0.48)}$ $71.96_{(0.49)}$
cutract (n=100) $72.36_{(0.47)}$ $70.96_{(0.68)}$ $70.20_{(0.45)}$ $\textbf{72.51}_{(0.55)}$ $69.97_{(0.97)}$ $71.94_{(0.89)}$ $67.94_{(1.01)}$ $72.42_{(0.66)}$ $70.53_{(1.51)}$ $71.98_{(1.39)}$ $55.72_{(2.04)}$ $69.14_{(0.93)}$ $67.59_{(0.71)}$ $72.42_{(0.54)}$ $68.79_{(0.83)}$ $69.68_{(0.76)}$ $66.89_{(1.03)}$ $71.52_{(0.70)}$
maggic (n=100) $67.46_{(0.07)}$ $59.65_{(0.50)}$ $58.98_{(0.23)}$ $\textbf{61.32}_{(0.42)}$ $55.71_{(0.83)}$ $58.90_{(0.72)}$ $57.20_{(0.91)}$ $59.34_{(0.64)}$ $57.26_{(0.50)}$ $58.28_{(0.46)}$ $49.54_{(0.71)}$ $57.91_{(0.74)}$ $56.36_{(0.54)}$ $60.11_{(0.54)}$ $58.89_{(0.51)}$ $58.99_{(0.42)}$ $56.17_{(0.68)}$ $58.86_{(0.69)}$
seer (n=100) $87.79_{(0.07)}$ $83.95_{(0.32)}$ $84.45_{(0.38)}$ $\textbf{85.37}_{(0.47)}$ $83.92_{(0.81)}$ $85.08_{(0.32)}$ $81.60_{(0.73)}$ $85.14_{(0.36)}$ $83.04_{(0.78)}$ $84.83_{(0.49)}$ $70.32_{(2.52)}$ $83.83_{(0.39)}$ $81.16_{(1.03)}$ $85.03_{(0.39)}$ $81.82_{(0.45)}$ $82.49_{(0.46)}$ $78.88_{(0.71)}$ $84.50_{(0.44)}$
compas (n=100) $67.18_{(0.30)}$ $62.56_{(0.72)}$ $68.02_{(0.29)}$ $\textbf{68.19}_{(0.37)}$ $60.10_{(1.60)}$ $62.47_{(1.11)}$ $60.01_{(1.16)}$ $63.73_{(0.99)}$ $58.32_{(0.91)}$ $61.34_{(0.94)}$ $59.97_{(0.82)}$ $64.19_{(0.83)}$ $60.02_{(0.63)}$ $64.04_{(0.60)}$ $61.44_{(0.84)}$ $61.73_{(0.81)}$ $59.97_{(1.00)}$ $62.82_{(1.05)}$
adult (n=100) $84.34_{(0.07)}$ $81.24_{(0.48)}$ $46.09_{(1.86)}$ $74.57_{(1.74)}$ $47.56_{(3.43)}$ $73.97_{(1.57)}$ $74.29_{(1.23)}$ $80.45_{(0.81)}$ $75.93_{(1.45)}$ $78.22_{(1.22)}$ $77.09_{(0.74)}$ $\textbf{81.66}_{(0.53)}$ $70.70_{(}{1.02)}$ $81.04_{(0.37)}$ $80.56_{(0.53)}$ $81.10_{(0.44)}$ $74.04_{(1.44)}$ $80.23_{(0.98)}$
drug (n=100) $78.00_{(0.19)}$ $73.58_{(0.72)}$ $76.24_{(0.64)}$ $\textbf{76.74}_{(0.51)}$ $69.46_{(2.43)}$ $71.05_{(2.30)}$ $68.19_{(2.06)}$ $73.28_{(1.41)}$ $72.43_{(0.96)}$ $73.79_{(0.65)}$ $67.26_{(1.22)}$ $75.28_{(0.46)}$ $62.67_{(2.23)}$ $73.12_{(0.97)}$ $70.90_{(1.08)}$ $71.53_{(1.13)}$ $68.22_{(1.60)}$ $73.59_{(0.97)}$
covid (n=200) $74.69_{(0.19)}$ $72.33_{(0.52)}$ $73.40_{(0.25)}$ $\textbf{74.62}_{(0.13)}$ $70.70_{(0.73)}$ $73.12_{(0.48)}$ $71.07_{(0.35)}$ $73.89_{(0.40)}$ $72.47_{(0.50)}$ $74.44_{(0.44)}$ $65.55_{(0.67)}$ $73.07_{(0.38)}$ $65.04_{(0.69)}$ $72.90_{(0.37)}$ $71.68_{(0.49)}$ $71.87_{(0.49)}$ $67.89_{(0.56)}$ $72.38_{(0.42)}$
cutract (n=200) $72.52_{(0.49)}$ $71.75_{(0.71)}$ $71.39_{(0.76)}$ $73.01_{(0.85)}$ $70.28_{(0.67)}$ $72.39_{(0.84)}$ $69.28_{(0.55)}$ $72.41_{(0.65)}$ $71.83_{(0.65)}$ $\textbf{74.03}_{(0.69)}$ $66.66_{(0.94)}$ $72.49_{(0.71)}$ $68.77_{(0.71)}$ $73.16_{(0.68)}$ $70.23_{(0.83)}$ $70.80_{(0.79)}$ $66.61_{(0.76)}$ $71.87_{(0.71)}$
maggic (n=200) $67.37_{(0.06)}$ $61.39_{(0.44)}$ $58.92_{(0.47)}$ $\textbf{61.41}_{(0.39)}$ $57.33_{(0.51)}$ $60.16_{(0.41)}$ $58.48_{(0.59)}$ $61.33_{(0.57)}$ $56.26_{(0.82)}$ $57.20_{(0.91)}$ $50.74_{(0.74)}$ $59.60_{(0.55)}$ $55.95_{(0.72)}$ $60.75_{(0.39)}$ $60.73_{(0.44)}$ $60.78_{(0.49)}$ $57.18_{(0.56)}$ $60.23_{(0.44)}$
seer (n=200) $87.84_{(0.08)}$ $84.63_{(0.35)}$ $84.39_{(0.19)}$ $85.56_{(0.30)}$ $83.48_{(0.41)}$ $84.80_{(0.34)}$ $82.04_{(0.70)}$ $85.34_{(0.36)}$ $84.39_{(0.44)}$ $\textbf{86.57}_{(0.27)}$ $82.15_{(0.43)}$ $86.03_{(0.23)}$ $77.73_{(1.65)}$ $85.19_{(0.22)}$ $83.38_{(0.32)}$ $84.15_{(0.42)}$ $79.71_{(0.58)}$ $85.26_{(0.25)}$
compas (n=200) $67.14_{(0.20)}$ $63.27_{(0.60)}$ $67.02_{(0.53)}$ $\textbf{68.15}_{(0.40)}$ $60.48_{(1.86)}$ $63.39_{(1.28)}$ $60.58_{(0.83)}$ $64.32_{(0.78)}$ $60.60_{(0.90)}$ $63.52_{(1.04)}$ $61.11_{(0.63)}$ $65.08_{(0.53)}$ $56.58_{(1.03)}$ $63.60_{(0.83)}$ $61.99_{(0.64)}$ $62.80_{(0.61)}$ $60.15_{(1.14)}$ $63.99_{(0.71)}$
adult (n=200) $84.25_{(0.04)}$ $82.12_{(0.41)}$ $40.96_{(2.47)}$ $75.84_{(1.43)}$ $49.89_{(3.22)}$ $76.11_{(1.43)}$ $78.18_{(0.26)}$ $82.32_{(0.29)}$ $81.66_{(0.19)}$ $83.17_{(0.20)}$ $80.06_{(0.53)}$ $\textbf{83.32}_{(0.35)}$ $74.31_{(0.81)}$ $82.64_{(0.33)}$ $82.26_{(0.38)}$ $82.39_{(0.33)}$ $75.21_{(0.69)}$ $82.02_{(0.25)}$
drug (n=200) $77.36_{(0.51)}$ $76.10_{(0.45)}$ $75.58_{(0.55)}$ $76.06_{(0.42)}$ $70.66_{(1.56)}$ $72.81_{(1.18)}$ $71.31_{(1.17)}$ $75.98_{(0.76)}$ $69.61_{(2.76)}$ $71.79_{(1.99)}$ $72.35_{(0.61)}$ $\textbf{77.41}_{(0.56)}$ $65.25_{(2.47)}$ $75.26_{(0.77)}$ $74.38_{(0.50)}$ $74.78_{(0.54)}$ $68.39_{(1.26)}$ $74.33_{(0.61)}$
C.3 Decoupling prior knowledge and data model
In-context samples Downstream accuracy
$n=1$ (Prior) $70.20\pm 1.60$
$n=20$ $73.87\pm 0.50$
$n=40$ $73.95\pm 0.67$
$n=100$ $74.71\pm 0.34$
$D_{\mathrm{oracle}}$ $74.6\pm 0.15$
[…]

### v3

@@ C.6 Full results for performance evaluation
seer (n=40) $87.92_{(0.08)}$ $82.93_{(0.55)}$ $84.29_{(0.39)}$ $\textbf{84.93}_{(0.46)}$ $83.46_{(1.20)}$ $84.44_{(0.69)}$ $80.05_{(0.93)}$ $83.67_{(0.51)}$ $82.59_{(1.48)}$ $81.37_{(1.11)}$ $54.93_{(2.06)}$ $81.11_{(1.24)}$ $79.88_{(0.53)}$ $84.36_{(0.62)}$ $80.79_{(0.67)}$ $82.21_{(0.67)}$ $78.69_{(2.32)}$ $83.62_{(1.06)}$
compas (n=40) $67.51_{(0.03)}$ $62.34_{(0.79)}$ $67.57_{(0.40)}$ $\textbf{67.85}_{(0.37)}$ $61.34_{(1.67)}$ $62.84_{(1.21)}$ $56.29_{(1.96)}$ $61.02_{(1.69)}$ $58.85_{(1.28)}$ $60.11_{(1.22)}$ $58.88_{(1.04)}$ $64.37_{(0.91)}$ $58.61_{(1.35)}$ $63.54_{(1.11)}$ $60.83_{(1.19)}$ $60.95_{(1.22)}$ $55.94_{(1.50)}$ $61.04_{(1.56)}$
adult (n=40) $84.17_{(0.10)}$ $79.44_{(1.03)}$ $48.31_{(3.58)}$ $73.82_{(2.19)}$ $49.21_{(1.78)}$ $74.27_{(1.49)}$ $71.82_{(1.22)}$ $79.11_{(0.71)}$ $71.51_{(1.85)}$ $77.99_{(0.65)}$ $66.77_{(3.32)}$ $78.81_{(1.36)}$ $71.13_{(2.27)}$ $79.71_{(1.02)}$ $77.90_{(1.07)}$ $78.84_{(1.19)}$ $72.58_{(1.34)}$ $\textbf{80.02}_{(0.76)}$
drug (n=40) $77.81_{(0.55)}$ $71.86_{(1.07)}$ $74.30_{(0.59)}$ $\textbf{75.79}_{(0.39)}$ $71.33_{(0.88)}$ $72.76_{(0.97)}$ $69.46_{(2.23)}$ $72.74_{(1.70)}$ $71.08_{(1.72)}$ $73.07_{(1.03)}$ $64.89_{(1.39)}$ $73.64_{(0.87)}$ $62.51_{(3.03)}$ $70.97_{(1.91)}$ $69.23_{(1.68)}$ $69.78_{(1.46)}$ $65.22_{(1.37)}$ $70.30_{(1.04)}$
covid (n=100) $74.41_{(0.11)}$ $71.57_{(0.48)}$ $73.77_{(0.27)}$ $\textbf{74.71}_{(0.34)}$ $70.71_{(0.46)}$ $72.76_{(0.44)}$ $69.05_{(0.96)}$ $72.13_{(0.66)}$ $71.60_{(0.59)}$ $73.22_{(0.46)}$ $63.52_{(1.29)}$ $72.04_{(0.57)}$ $64.25_{(1.51)}$ $72.64_{(0.64)}$ $70.08_{(0.67)}$ $70.78_{(0.59)}$ $69.05_{(0.48)}$ $71.96_{(0.49)}$
cutract (n=100) $72.23_{(0.65)}$ $70.96_{(0.68)}$ $70.20_{(0.45)}$ $\textbf{72.51}_{(0.55)}$ $69.97_{(0.97)}$ $71.94_{(0.89)}$ $67.94_{(1.01)}$ $72.42_{(0.66)}$ $70.53_{(1.51)}$ $71.98_{(1.39)}$ $55.72_{(2.04)}$ $69.14_{(0.93)}$ $67.59_{(0.71)}$ $72.42_{(0.54)}$ $68.79_{(0.83)}$ $69.68_{(0.76)}$ $66.89_{(1.03)}$ $71.52_{(0.70)}$
maggic (n=100) $67.50_{(0.04)}$ $59.65_{(0.50)}$ $58.98_{(0.23)}$ $\textbf{61.32}_{(0.42)}$ $55.71_{(0.83)}$ $58.90_{(0.72)}$ $57.20_{(0.91)}$ $59.34_{(0.64)}$ $57.26_{(0.50)}$ $58.28_{(0.46)}$ $49.54_{(0.71)}$ $57.91_{(0.74)}$ $56.36_{(0.54)}$ $60.11_{(0.54)}$ $58.89_{(0.51)}$ $58.99_{(0.42)}$ $56.17_{(0.68)}$ $58.86_{(0.69)}$
seer (n=100) $87.92_{(0.08)}$ $83.95_{(0.32)}$ $84.45_{(0.38)}$ $\textbf{85.37}_{(0.47)}$ $83.92_{(0.81)}$ $85.08_{(0.32)}$ $81.60_{(0.73)}$ $85.14_{(0.36)}$ $83.04_{(0.78)}$ $84.83_{(0.49)}$ $70.32_{(2.52)}$ $83.83_{(0.39)}$ $81.16_{(1.03)}$ $85.03_{(0.39)}$ $81.82_{(0.45)}$ $82.49_{(0.46)}$ $78.88_{(0.71)}$ $84.50_{(0.44)}$
compas (n=100) $67.51_{(0.03)}$ $62.56_{(0.72)}$ $68.02_{(0.29)}$ $\textbf{68.19}_{(0.37)}$ $60.10_{(1.60)}$ $62.47_{(1.11)}$ $60.01_{(1.16)}$ $63.73_{(0.99)}$ $58.32_{(0.91)}$ $61.34_{(0.94)}$ $59.97_{(0.82)}$ $64.19_{(0.83)}$ $60.02_{(0.63)}$ $64.04_{(0.60)}$ $61.44_{(0.84)}$ $61.73_{(0.81)}$ $59.97_{(1.00)}$ $62.82_{(1.05)}$
adult (n=100) $84.17_{(0.10)}$ $81.24_{(0.48)}$ $46.09_{(1.86)}$ $74.57_{(1.74)}$ $47.56_{(3.43)}$ $73.97_{(1.57)}$ $74.29_{(1.23)}$ $80.45_{(0.81)}$ $75.93_{(1.45)}$ $78.22_{(1.22)}$ $77.09_{(0.74)}$ $\textbf{81.66}_{(0.53)}$ $70.70_{(}{1.02)}$ $81.04_{(0.37)}$ $80.56_{(0.53)}$ $81.10_{(0.44)}$ $74.04_{(1.44)}$ $80.23_{(0.98)}$
drug (n=100) $77.81_{(0.55)}$ $73.58_{(0.72)}$ $76.24_{(0.64)}$ $\textbf{76.74}_{(0.51)}$ $69.46_{(2.43)}$ $71.05_{(2.30)}$ $68.19_{(2.06)}$ $73.28_{(1.41)}$ $72.43_{(0.96)}$ $73.79_{(0.65)}$ $67.26_{(1.22)}$ $75.28_{(0.46)}$ $62.67_{(2.23)}$ $73.12_{(0.97)}$ $70.90_{(1.08)}$ $71.53_{(1.13)}$ $68.22_{(1.60)}$ $73.59_{(0.97)}$
covid (n=200) $74.41_{(0.11)}$ $72.33_{(0.52)}$ $73.40_{(0.25)}$ $\textbf{74.62}_{(0.13)}$ $70.70_{(0.73)}$ $73.12_{(0.48)}$ $71.07_{(0.35)}$ $73.89_{(0.40)}$ $72.47_{(0.50)}$ $74.44_{(0.44)}$ $65.55_{(0.67)}$ $73.07_{(0.38)}$ $65.04_{(0.69)}$ $72.90_{(0.37)}$ $71.68_{(0.49)}$ $71.87_{(0.49)}$ $67.89_{(0.56)}$ $72.38_{(0.42)}$
cutract (n=200) $72.23_{(0.65)}$ $71.75_{(0.71)}$ $71.39_{(0.76)}$ $73.01_{(0.85)}$ $70.28_{(0.67)}$ $72.39_{(0.84)}$ $69.28_{(0.55)}$ $72.41_{(0.65)}$ $71.83_{(0.65)}$ $\textbf{74.03}_{(0.69)}$ $66.66_{(0.94)}$ $72.49_{(0.71)}$ $68.77_{(0.71)}$ $73.16_{(0.68)}$ $70.23_{(0.83)}$ $70.80_{(0.79)}$ $66.61_{(0.76)}$ $71.87_{(0.71)}$
maggic (n=200) $67.50_{(0.04)}$ $61.39_{(0.44)}$ $58.92_{(0.47)}$ $\textbf{61.41}_{(0.39)}$ $57.33_{(0.51)}$ $60.16_{(0.41)}$ $58.48_{(0.59)}$ $61.33_{(0.57)}$ $56.26_{(0.82)}$ $57.20_{(0.91)}$ $50.74_{(0.74)}$ $59.60_{(0.55)}$ $55.95_{(0.72)}$ $60.75_{(0.39)}$ $60.73_{(0.44)}$ $60.78_{(0.49)}$ $57.18_{(0.56)}$ $60.23_{(0.44)}$
seer (n=200) $87.92_{(0.08)}$ $84.63_{(0.35)}$ $84.39_{(0.19)}$ $85.56_{(0.30)}$ $83.48_{(0.41)}$ $84.80_{(0.34)}$ $82.04_{(0.70)}$ $85.34_{(0.36)}$ $84.39_{(0.44)}$ $\textbf{86.57}_{(0.27)}$ $82.15_{(0.43)}$ $86.03_{(0.23)}$ $77.73_{(1.65)}$ $85.19_{(0.22)}$ $83.38_{(0.32)}$ $84.15_{(0.42)}$ $79.71_{(0.58)}$ $85.26_{(0.25)}$
compas (n=200) $67.51_{(0.03)}$ $63.27_{(0.60)}$ $67.02_{(0.53)}$ $\textbf{68.15}_{(0.40)}$ $60.48_{(1.86)}$ $63.39_{(1.28)}$ $60.58_{(0.83)}$ $64.32_{(0.78)}$ $60.60_{(0.90)}$ $63.52_{(1.04)}$ $61.11_{(0.63)}$ $65.08_{(0.53)}$ $56.58_{(1.03)}$ $63.60_{(0.83)}$ $61.99_{(0.64)}$ $62.80_{(0.61)}$ $60.15_{(1.14)}$ $63.99_{(0.71)}$
adult (n=200) $84.17_{(0.10)}$ $82.12_{(0.41)}$ $40.96_{(2.47)}$ $75.84_{(1.43)}$ $49.89_{(3.22)}$ $76.11_{(1.43)}$ $78.18_{(0.26)}$ $82.32_{(0.29)}$ $81.66_{(0.19)}$ $83.17_{(0.20)}$ $80.06_{(0.53)}$ $\textbf{83.32}_{(0.35)}$ $74.31_{(0.81)}$ $82.64_{(0.33)}$ $82.26_{(0.38)}$ $82.39_{(0.33)}$ $75.21_{(0.69)}$ $82.02_{(0.25)}$
drug (n=200) $77.81_{(0.55)}$ $76.10_{(0.45)}$ $75.58_{(0.55)}$ $76.06_{(0.42)}$ $70.66_{(1.56)}$ $72.81_{(1.18)}$ $71.31_{(1.17)}$ $75.98_{(0.76)}$ $69.61_{(2.76)}$ $71.79_{(1.99)}$ $72.35_{(0.61)}$ $\textbf{77.41}_{(0.56)}$ $65.25_{(2.47)}$ $75.26_{(0.77)}$ $74.38_{(0.50)}$ $74.78_{(0.54)}$ $68.39_{(1.26)}$ $74.33_{(0.61)}$
C.7 CLLM  with context in low-resource languages
We investigate how the language of tabular column names affect the performance of CLLM. To do so, for a given dataset, we translate the feature names from English to 2 low-resource languages (Ranathunga et al., 2023), i.e. to (i) Swahili and (ii) Hausa. Changing the language of the features alters the contextual information provided to the LLM to generate synthetic data.
We report the results in Table 14. As expected, we observe for both Swahili and Hausa a performance drop compared to using English feature names, thereby mirroring the results in Section 2.1 on the importance of contextual information in the prompt. However, we note that CLLM  for both Swahili and Hausa remains highly competitive compared to other baselines, notably thanks to the curation mechanism.
Table 14: Test AUC in settings where the names of features have been translated to Swahili and Hausa, with and without curation
Dataset Uncurated English Curated English Uncurated Swahili Curated Swahili Uncurated Hausa Curated Hausa
covid ($n=20$) 73.78 73.87 71.32 73.06 68.75 70.01
seer ($n=20$) 84.53 84.82 83.92 85.47 81.57 82.73
compas ($n=20$) 68.01 67.91 64.42 61.61 62.85 63.06
covid $n=40$ 73.40 73.95 71.53 73.20 70.15 72.13
seer $n=40$ 84.29 84.93 79.71 84.25 81.47 84.66
compas $n=40$ 67.57 67.85 64.92 65.78 62.31 63.20
Takeaways:
  1. 1.
    CLLM  can provide valuable augmentation and performance gains even with feature names in low-resource languages.
  2. 2.
    For optimal performance, we recommend using CLLM  with English feature names or translating the feature names to English before applying CLLM .
C.8 Comparison of curation mechanism vs Data-IQ
We compare CLLM’s curation mechanism with the approach taken by Data-IQ (Seedat et al., 2022a).
They differ along four dimensions:
  1. 1.
    Problem setting (data augmentation vs data understanding): CLLM  focuses on the problem of data augmentation in low-data regimes by curating synthetic samples. In contrast, Data-IQ aims to understand and characterize subgroups within a given real dataset.
  2. 2.
    Conceptually: CLLM  curates a large synthetic dataset $D_{\mathrm{syn}}$ with respect to a very small gold standard real dataset $D_{\mathrm{train}}$, whereas Data-IQ aims to characterize subgroups of learnable samples in a single large real dataset $D$ (e.g. to find the hard samples in $D$).
  3. 3.
    Technically: CLLM  trains the curation model only on the small but gold standard $D_{\mathrm{train}}$. The learning dynamics of the synthetic samples in $D_{\mathrm{syn}}$ are then assessed with respect to this curation model. In contrast, Data-IQ computes learning dynamics for real samples in $D$ using a curator model trained on the same $D$ it assesses.
  4. 4.
    Empirical performance: In CLLM , the curation aims at discarding synthetic samples which contradict the learning signal obtained from the real data. If we were to perform the curation like in Data-IQ, this implies that we would instead have to merge $D_{\mathrm{train}}$ and $D_{\mathrm{syn}}$, and train a curator model on $D_{\mathrm{merged}}=D_{\mathrm{train}}\cup D_{\mathrm{syn}}$ to assess $D_{\mathrm{syn}}$. Intuitively, with such an approach, the signal from the small $D_{\mathrm{train}}$ would be overshadowed by $D_{\mathrm{syn}}$, as the latter is a lot bigger in size ($1000$ samples), thus making the curation irrelevant.
We now show empirical evidence regarding the last point, by comparing CLLM  with a baseline where we train the curator on $D_{\mathrm{merged}}=D_{\mathrm{train}}\cup D_{\mathrm{syn}}$. The results are in Table 15. We observe that:
  1. 1.
    CLLM’s curation outperforms Data-IQ curation on downstream performance across all datasets and $n$.
  2. 2.
    Data-IQ curation does not always improve upon the uncurated baseline, which aligns with our intuition that $D_{\mathrm{syn}}$ overshadows $D_{\mathrm{train}}$ because of its size.
Table 15: Comparing the CLLM  curation mechanism to Data-IQ curation (Seedat et al., 2022a). We report the AUC averaged over 4 downstream models on $\mathcal{D}_{\mathrm{test}}$.
Dataset Uncurated Curated (CLLM ) - Ours Curated (Data-IQ)
covid (n=20) 73.78 73.87 73.45
cutract (n=20) 71.15 72.50 70.84
maggic (n=20) 60.70 61.48 60.68
seer (n=20) 84.53 84.82 83.97
compas (n=20) 68.01 67.91 67.45
adult (n=20) 50.39 71.48 43.59
drug (n=20) 75.08 75.29 74.23
covid (n=40) 73.40 73.95 73.36
cutract (n=40) 69.87 71.72 69.57
maggic (n=40) 59.29 60.77 59.65
seer (n=40) 84.29 84.93 84.47
compas (n=40) 67.57 67.85 67.07
adult (n=40) 48.31 73.82 48.92
drug (n=40) 74.30 75.79 74.45
covid (n=100) 73.77 74.71 73.77
cutract (n=100) 70.20 72.51 70.29
maggic (n=100) 58.98 61.32 58.66
seer (n=100) 84.45 85.37 84.33
compas (n=100) 68.02 68.19 67.82
adult (n=100) 46.09 74.57 45.13
drug (n=100) 76.24 76.74 76.11
covid (n=200) 73.40 74.62 73.27
cutract (n=200) 71.39 73.01 71.43
maggic (n=200) 58.92 61.41 58.71
seer (n=200) 84.39 85.56 84.49
compas (n=200) 67.02 68.15 67.26
adult (n=200) 40.96 75.84 41.01
drug (n=200) 75.58 76.06 75.39
C.9 Comparing CLLM  vs TabPFN
We outlined in Appendix A various dimensions on which CLLM  and transfer learning / meta-learning / few-shot learning differ. To provide further empirical evidence on why the above dimensions are important (notably the point on the choice of downstream backbone), we compare CLLM  with TabPFN (Hollmann et al., 2022), a few-shot learning method designed for small tabular problems. We chose TabPFN because it meets our criteria where it does not require access to external datasets, since the model is pretrained on an extensive set of synthetic tabular datasets and can perform few-shot learning with its transformer backbone.
We use the pretrained model released by the authors at https://github.com/automl/TabPFN and show the results in Table 16.
Table 16: Comparison of CLLM  vs TabPFN. We report the AUC averaged over 4 downstream models on $\mathcal{D}_{\mathrm{test}}$.
Dataset Uncurated Curated (CLLM ) TabPFN
covid (n=20) 73.78 73.87 66.31
cutract (n=20) 71.15 72.50 63.86
maggic (n=20) 60.70 61.48 55.49
seer (n=20) 84.53 84.82 75.30
compas (n=20) 68.01 67.91 57.04
adult (n=20) 50.39 71.48 71.70
drug (n=20) 75.08 75.29 69.18
covid (n=40) 73.40 73.95 67.26
cutract (n=40) 69.87 71.72 65.93
maggic (n=40) 59.29 60.77 57.75
seer (n=40) 84.29 84.93 79.48
compas (n=40) 67.57 67.85 57.72
adult (n=40) 48.31 73.82 75.36
drug (n=40) 74.30 75.79 71.32
covid (n=100) 73.77 74.71 69.50
cutract (n=100) 70.20 72.51 70.01
maggic (n=100) 58.98 61.32 59.07
seer (n=100) 84.45 85.37 80.96
compas (n=100) 68.02 68.19 63.42
adult (n=100) 46.09 74.57 77.36
drug (n=100) 76.24 76.74 72.68
covid (n=200) 73.40 74.62 70.82
cutract (n=200) 71.39 73.01 71.59
maggic (n=200) 58.92 61.41 60.67
seer (n=200) 84.39 85.56 82.05
compas (n=200) 67.02 68.15 65.51
adult (n=200) 40.96 75.84 79.18
drug (n=200) 75.58 76.06 74.29
Takeaways: We see CLLM  outperforms TabPFN on $6/7$ datasets, for the different $n$. The performance gains are especially noticeable in the ultra low-sample regime ($n=20$).
Finally, we acknowledge that CLLM  is not a one-size-fits-all approach. When the assumptions underpinning transfer learning or meta-learning hold (e.g. availability of external datasets), combining ideas from these learning paradigms with the augmentation methodology of CLLM  could constitute an interesting research direction, but this falls beyond the scope of our current work.
C.10 CLLM  in specialized domains.
We conduct an experiment with the additional dataset Higgs (Whiteson, 2014). We chose this dataset because it comes from a specialized domain (physics), where the dataset consists in kinematic properties of particles measured by the particle detectors of an accelerator, which is likely under-represented in the LLMs training corpus.
We note that the contextual information for this dataset is quite specific, as can be seen from the names of the features, which include for example ”m_bb”, ”m_jj”, ”m_jjj”. This particular contextual information, which is not as semantically meaningful as other datasets, makes it interesting to compare the performance of CLLM  with the traditional baselines. We show the results in Table 17. As we can see, the performance of CLLM  is good for the smaller $n$. As $n$ increases, the downstream task benefits more from the use of the other baselines. We highlight that in this case, our curation mechanism still benefits both CLLM  and the baselines.
Table 17: CLLM  performance on the Higgs dataset
Dataset CLLM CTGAN TVAE NFLOW
Uncur Cur Uncur Cur Uncur Cur Uncur Cur
higgs (n=20) 65.82 70.25 62.73 67.31 59.14 69.89 56.45 69.63
higgs (n=40) 65.38 71.01 63.07 70.41 59.29 67.61 60.27 71.54
higgs (n=100) 68.62 73.42 68.46 75.59 68.90 76.05 59.73 75.86
higgs (n=200) 66.39 75.18 74.27 79.75 73.65 79.32 60.99 79.42
C.11 Behavior beyond the low-data regime.
We examine the behavior of baselines beyond $n=200$ and the low-data regime. We provide a plot in Figure 10 for three datasets examining the behavior of CTGAN, TVAE and NFLOW for increasing $n$. We fix the CLLM  method at $n=200$ (as we do not see too much of an additional performance increase, due to the LLM context window which limits the number of in-context samples we can provide).
Figure 10: Behavior of baselines at high $n$, i.e. beyond the low data regime.
Takeaways:
  1. 1.
    Other baselines initially improve with more real samples, however their performance gains tend to plateau out at around $n\in\{200,400\}$ samples. This suggests we only get minimal gains as increasing numbers of samples are added to the baselines.
  2. 2.
    CLLM  either outperforms the baselines even for very high $n$ (Compas, Covid) or remains competitive (SEER), even when the number of samples used by CTGAN, TVAE and NFLOW is twice or thrice bigger.
C.12 Addressing potential biases from the LLM
There are many challenges which arise from the use of LLMs. An example is potential biases from the LLM which might stem from their training data and might affect our synthetic data generated by the LLM.
We first describe some considerations which show how a practicioner using CLLM  can address this issue.
  1. 1.
    Choice of LLM: Different LLMs may exhibit different levels of bias, due to their respective pretraining and alignment with human feedback. With CLLM , the practitioner has a lot of flexibility in the choice of the LLM. This is an important aspect when it comes to bias. The practitioner may want to minimize the risk of having bias in the augmented data by selecting one particular LLM for their task based on prior knowledge about different LLMs and their respective potential bias (Gallegos et al., 2023).
  2. 2.
    Fairness analysis on the augmented data: Practitioners can directly assess potential bias in the data generated by CLLM : (i) by doing a standard fairness analysis and computing different fairness metrics (Dwork et al., 2012; Hardt et al., 2016; Mehrabi et al., 2021) on the generated data, as can be done more generally when building a model with any given dataset. (ii) If some bias is detected, the practitioner can use any off-the-shelf debiasing method (Calmon et al., 2017; Feldman et al., 2015), or use a training objective to make the downstream model fair (Berk et al., 2017; Zhang et al., 2018). Note that this approach is possible because CLLM  is a data augmentation method, which gives a lot of flexibility for the practitioner to adjust the data according to their needs.
  3. 3.
    CLLM can reduce representation bias: While these aforementioned considerations involve choices external to CLLM , we also want to emphasize that our CLLM  approach has the potential to already address bias indirectly. As shown in Figure 3, underrepresented groups in the population benefit the most from CLLM . This suggests that we could use CLLM  to remove representation bias, by leveraging CLLM  to augment $\mathcal{D}_{\mathrm{train}}$ with synthetic data from underrepresented groups. Table 1 and Table 2 show that data generated with CLLM  better aligns with the ground-truth distribution, compared to using a widely used conventional generative approach (TVAE).
  4. 4.
[…]